OpenStudy (anonymous):

Find the equation of the line through (-4, 6) and perpendicular to y = -3x + 4

OpenStudy (anonymous):

slope of y=-3x + 4 is equal to -3 therefore slope of desired line is 1/3 (m1 * m2 = -1 for perpendicular lines) therefore the equation of desired line is y-(-4) = 1/3(x-6) => y+4 = 1/3(x-6) => 3y + 12 = x- 6 => x-3y = 18